please help, will rate thanks

Transverse waves traveling along a string have the following properties.

Amplitude of the wave = 2.2 mm
Wavelength of the wave = 0.15 m
Speed of the wave = 343 m/s

Determine the time for a particle of the string to move through a total distance of 1.5 km.
1 s
Answer #1

The frequency = v / lambda = 343 / 0.15

= 2286.7 Hz

The particle moves four times the amplitude in one period

average transverse velocity v = (4 * 2.2 * 10-3) * 2286.7

= 20.12 m/s

It will travel 1.5 km in t = d/v = 1500 / 20.12

= 74.54 s.
Answer #2


T = wavelength/v = 0.15/343 = 4.373e-4


in 1 T the particle travells a distance of 4A = 8.8 mm

time taken to travel a distance of 1.5 km


t = (4.373e-4*1500)/8.8e-3 = 74.54s
Answer #3

u10l2e2.gif

Since the period is the reciprocal of the frequency, the expression 1/f can be substituted into the above equation for period.

Speed = Wavelength